# Quantum Non-Local Nonstabilizerness

### v3

@@ Abstract
NN in many-body state. Building on our improved understanding of the NN in the two-qubit case, we now shift our focus to the scaling behavior of the two-point NN in many-body quantum states, defined as $\mathcal{M}(\rho_{1,r})$, where $\rho_{1,r}$ represents the reduced density matrix for the first qubit and the $r$-th qubit. For systems with translational symmetry and periodic boundary conditions, the two-point NN depends solely on the relative separation $r$. The scaling behavior provides valuable information on the properties of many-body states, especially for critical states. We also explore the measurement-induced non-local nonstabilizerness (MINN), defined as ~$\widetilde{\mathcal{M}}(r)\equiv\sum_{i}p_{i}\mathcal{M}(\left|\Psi_{1,r}^{i}\right\rangle)$, where $p_{i}$ is the probability of obtaining the outcome $i$ from projective measurements performed on all qubits except the first and the $r$-th qubits, and $\left|\Psi_{1,r}^{i}\right\rangle$ represents the corresponding post-measurement state. Numerical details are provided in Supplemental Material [47].
@@
We first consider the TFIM, with Hamiltonian given by $H_{\text{TFIM}}=-\sum_{i}\sigma_{x}^{i}\sigma_{x}^{i+1}-h\sum_{i}\sigma_{z}^{i}$. The ground state is paramagnetic for $h>1$ and ferromagnetic for $h<1$, with a quantum phase transition occurring at $h=1$. Due to the $\mathbb{Z}_{2}$ symmetry $\prod_{i}\sigma_{x}^{i}$, the Pauli basis representation of $\rho_{1,r}$ is constrained to be in the canonical form, enabling direct computation of the NN. As shown in Fig. 2(a), the NN exhibits a power-law decay with an exponent of around $0.5$, close to that of mutual information. In [47], we show that their decay exponents are analytically identical for sufficiently large $r$. Away from criticality, the NN decays exponentially to a constant value, distinguishing it from critical states.
Next, we examine the MINN for different measurement directions. As illustrated in Fig. 2(b-d), the MINN exhibits a power-law decay for critical states, with different decay exponents depending on the measurement axis. The MINN under $y-$ and $z-$axes measurements decays more slowly than the pre-measurement behavior, but still faster than the slowest decaying correlation prior to measurement, namely the $\left\langle\sigma_{x}^{1}\sigma_{x}^{r}\right\rangle$, which decays with exponent $1/4$ [73]. For non-critical states, the MINN either becomes constant or decays exponentially.
@@
Figure 3: Results for critical MHC with $L=48$, with a maximum bond dimension $\chi=1000$. (a) Haar-random unitaries are applied in a brick-wall manner, with a measurement rate of $p=0.17$ [43]. For the MINN, all qubits except the first and the $r$-th qubits are projectively measured along the $z$-axis. (b) A comparison between averaged two-point NN $\overline{\mathcal{M}}$ and total nonstabilizerness $\overline{M}$. Each data point is averaged over 3000 trajectories. The inset shows the power-law fit with a constant shift for $\overline{\mathcal{M}}$. (c) The MINN and the average total nonstabilizerness after measurement, with each data point averaged over $4\times 10^{4}$ trajectories. (d) A comparison between the scaling behavior of the MINN and the mutual information $I_{E}$ prior measurement. The decay exponents for $\widetilde{\mathcal{M}}$ and $I_{E}$ are around $0.76$ and $3.31$, respectively.
@@
We then consider the MHC at a critical measurement rate of $p=0.17$. The circuit structure is shown in Fig. 3(a). We consider the average NN defined as ¯$\overline{\mathcal{M}}=\mathbb{E}_{U}\mathbb{E}_{m}[\mathcal{M}]$, where $\mathbb{E}_{U}$ and $\mathbb{E}_{m}$ represent the averaging over unitary gates and the measurement outcomes, respectively. Since the model is structureless, with only locality as a constraint, $\rho_{1,r}$ does not necessarily have a canonical Pauli basis representation, and the NN must be computed by numerical optimization. As shown in Fig. 3(b), NN is significantly lower than total nonstabilizerness and also exhibits a power-law decay. We then consider the MINN and compare it with the average total nonstabilizerness for post-measurement states, as presented in Fig. 3(c). The measurement direction is insignificant, and we assume it to be along the $z$-axis. It is evident that a greater amount of nonstabilizerness can be removed by local unitary transformations as the distance increases. Furthermore, we observe that the decay exponent of MINN is less than half of that of mutual information prior to measurement, as shown in Fig. 3(d). Since the slowest correlation a state can maintain must decay at an exponent at least half that of mutual information [74], this suggests that the measurement induces NN with a longer range than any pre-existing correlations. This phenomenon is akin to entanglement swapping, where measurements induce long-range entanglement between previously uncorrelated qubits. We refer to this effect as “nonstabilizerness swapping”. Since it has been established that $\mathcal{M}$ is upper bounded by entanglement [38], and that measurement-induced entanglement cannot decay more slowly than correlations in sign-free states [75], it follows that nonstabilizerness swapping can only occur in states possessing sign structures.
Discussions. As we focus on the NN for qubit systems, a natural extension is to investigate systems with higher-dimensional local Hilbert spaces, such as qutrits. In systems where the local Hilbert space dimension is an odd prime, mana serves as an alternative, appropriate nonstabilizerness measure for both pure and mixed states [76, 77, 78, 79]. Notably, like SRE, mana does not necessitate any minimization procedure. However, it remains an open question whether an analytical expression can be derived for the two-qutrit case. It would also be interesting to investigate the consistency between the results obtained using SRE and mana. We leave this to future work.
While NN can be expressed analytically for two-qubit pure states, making it computationally straightforward and feasible for experimental measurements, the situation becomes more complex for general multi-qubit states. In this case, computing the NN presents an optimization problem involving $O(N^{2})$ parameters, which becomes increasingly difficult as the number of qubits grows. An approach to address this challenge is to employ automatic differentiation techniques, similar to those used in simulating the variational quantum eigensolver algorithm, with the target function being the total nonstabilizerness rather than the energy [80]. Scaling up the calculation of NN would facilitate the classification of states based on whether they exhibit volume-law or area-law NN, analogous to the classification of states in terms of entanglement. We conjecture that only states that exhibit both volume-law entanglement entropy and volume-law nonstabilizerness can sustain a volume-law NN.
@@
Additionally, it is noteworthy that a related measure, nonstabilizerness entanglement entropy (NsEE), has recently been introduced [81, 82]. In this formulation, free operations are defined as global Clifford gates instead of local unitary gates, and the computed quantity is the entanglement entropy of the total state rather than its nonstabilizerness. While NsEE and NN aim to characterize the “quantumness” of a system from distinct perspectives, further investigation into the relationship between these two measures would be a valuable direction for future research.
@@ Acknowledgements.
Acknowledgments. We thank Xiaoyu Dong for valuable discussions. This work is supported by the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600, No. 24LZ1400100 and No. 2019SHZDZX01.
Supplemental Material for “Quantum Non-Local Nonstabilizerness”
Dongheng Qian1, 2 and Jing Wang1, 2, 3, ${}^{1,\,2,\,3,\,4,\,*}$
1State Key Laboratory of Surface Physics and Department of Physics, Fudan University, Shanghai 200433, China
2Shanghai Research Center for Quantum Sciences, Shanghai 201315, China
3Institute for Nanoelectronic Devices and Quantum Computing, Fudan University, Shanghai 200433, China
4Hefei National Laboratory, Hefei 230088, China
S1 Analytic results
S1.1 Proof of sub-additivity
In this section, we prove that the non-local nonstabilizerness (NN), defined as $\mathcal{M}_{AB}(\left|\Psi_{AB}\right\rangle)=\min_{U_{A}\otimes U_{B}}M(U_{A}\otimes U_{B}\left|\Psi_{AB}\right\rangle)$, is sub-additive if the nonstabilizerness measure $M$ is additive. By definition, we have:
$\mathcal{M}_{AB}(\left|\Psi_{A_{1}B_{1}}\right\rangle\otimes\left|\Phi_{A_{2}B_{2}}\right\rangle)=\min_{{U_{A}\otimes U_{B}}}M(\left|\Psi_{A_{1}B_{1}}\right\rangle\otimes\left|\Phi_{A_{2}B_{2}}\right\rangle)=\min_{U_{A}\otimes U_{B}}[M(\left|\Psi_{A_{1}B_{1}}\right\rangle)+M(\left|\Phi_{A_{2}B_{2}}\right\rangle)].$ (S1)
@@
$\mathcal{M}_{AB}(\left|\Psi_{A_{1}B_{1}}\right\rangle\otimes\left|\Phi_{A_{2}B_{2}}\right\rangle)\leq\min_{U_{A_{1}}\otimes U_{B_{1}}}M(\left|\Psi_{A_{1}B_{1}}\right\rangle)+\min_{U_{A_{2}}\otimes U_{B_{2}}}M(\left|\Phi_{A_{2}B_{2}}\right\rangle)=\mathcal{M}_{A_{1}B_{1}}(\left|\Psi_{A_{1}B_{1}}\right\rangle)+\mathcal{M}_{A_{2}B_{2}}(\left|\Phi_{A_{2}B_{2}}\right\rangle).$ (S2)
@@
Building on this sub-additivity, it is straightforward to establish an upper bound for the NN of a many-body state, provided that it can be brought into product states of two-qubit pure states in canonical form via local unitary transformations. It has been shown that such transformations can be approximately realized using the Petz recovery map for Haar-random states [1]. However, the exact determination of such a transformation remains an open problem and is a subject worth further exploration.
S1.2 Proof of theorem
@@
$\rho=\frac{1}{4}\sum_{\mu,\nu}R_{\mu\nu}\sigma_{\mu}\sigma_{\nu},$ (S3)
where $\mu,\nu\in\{0,x,y,z\}$, and the coefficients $R$ are determined by $R_{\mu\nu}=\text{Tr}(\rho\sigma_{\mu}\otimes\sigma_{\nu})$ [2, 3]. The matrix $R$ takes a block structure of the form:
$R=\begin{pmatrix}1&\textbf{a}^{T}\\ \textbf{b}&T\end{pmatrix}$ (S4)
where a and b are the Bloch vectors corresponding to the reduced density matrices of qubit $A$ and $B$, respectively, and $T$ represents the correlation matrix. In this representation, the second stabilizer Rényi entropy (SRE) of the state $\rho$ is concisely expressed as:
$M(\rho)=-\text{log}\left(\sum_{\mu\nu}{|R_{\mu\nu}|^{4}}\right)-S_{2}(\rho),$ (S5)
@@
$\sum_{i}\left|a_{i}\right|^{4}\leq\sum_{i}\left|a_{i}\right|^{2}=\text{const},$ (S6)
@@
$\sum_{i=1}^{3}\left|T^{\prime}_{i,j}\right|^{4}=\sum_{i=1}^{3}\left|d_{j}O_{A,ij}\right|^{4}=\left|d_{j}\right|^{4}\sum_{i=1}^{3}\left|O_{A,ij}\right|^{4}\leq\left|d_{j}\right|^{4}\sum_{i=1}^{3}\left|O_{A,ij}\right|^{2}=\left|d_{j}\right|^{4},$ (S7)
@@
$R=\begin{pmatrix}1&0&0&\text{cos}(2\theta)\\ 0&\text{sin}(2\theta)&0&0\\ 0&0&-\text{sin}(2\theta)&0\\ \text{cos}(2\theta)&0&0&1\end{pmatrix},$ (S8)
@@
It is worth noting that while a local unitary transformation always exists that brings a two-qubit pure state into canonical form, this is not generally true for a mixed state. However, a unique normal form exists for any two-qubit mixed state, where the correlation matrix $T$ is diagonal, although the Bloch vectors of the qubits need not be fully polarized [4].
S1.3 Probability distribution
In this section, we derive the analytic expression for the probability distribution of NN for the two-qubit Haar random state. It is well known that the Haar random state has an entanglement spectrum $\{\text{cos}(\theta),\text{sin}(\theta)\}$, with the associated probability distribution given by $P(x)=3(1-2x)^{2}$, where we define $x\equiv\text{cos}^{2}(\theta)$ [5, 6]. From the corollary stated in the main text, the NN for a state with such an entanglement spectrum is expressed as:
$\mathcal{M}(x)=\text{log}\left(\frac{8}{7+\text{cos}(8\theta)}\right)=\text{log}\left(\frac{1}{1+4x(x-1)(1-2x)^{2}}\right).$ (S9)

### v4

@@ V NN in many-body states
Building on our improved understanding of the NN in the two-qubit case, we now shift our focus to the scaling behavior of the two-point NN in many-body quantum states, defined as $\mathcal{M}(\rho_{1,r})$, where $\rho_{1,r}$ represents the reduced density matrix for the first qubit and the $r$-th qubit. For systems with translational symmetry and periodic boundary conditions, the two-point NN depends solely on the relative separation $r$. The scaling behavior provides valuable information on the properties of many-body states, especially for critical states. We also explore the measurement-induced non-local nonstabilizerness (MINN), defined as ~$\widetilde{\mathcal{M}}(r)\equiv\sum_{i}p_{i}\mathcal{M}(\left|\Psi_{1,r}^{i}\right\rangle)$, where $p_{i}$ is the probability of obtaining the outcome $i$ from projective measurements performed on all qubits except the first and the $r$-th qubits, and $\left|\Psi_{1,r}^{i}\right\rangle$ represents the corresponding post-measurement state. Numerical details are provided in Appendix E.
@@
We first consider the TFIM, with Hamiltonian given by $H_{\text{TFIM}}=-\sum_{i}\sigma_{x}^{i}\sigma_{x}^{i+1}-h\sum_{i}\sigma_{z}^{i}$. The ground state is paramagnetic for $h>1$ and ferromagnetic for $h<1$, with a quantum phase transition occurring at $h=1$. Due to the $\mathbb{Z}_{2}$ symmetry $\prod_{i}\sigma_{x}^{i}$, the Pauli basis representation of $\rho_{1,r}$ is constrained to be in the canonical form, enabling direct computation of the NN. As shown in Fig. 2(a), the NN exhibits a power-law decay with an exponent of around $0.5$, close to that of mutual information. In Appendix D, we show that their decay exponents are analytically identical for sufficiently large $r$. Away from criticality, the NN decays exponentially to a constant value, distinguishing it from critical states. Next, we examine the MINN for different measurement directions. As illustrated in Fig. 2(b-d), the MINN exhibits a power-law decay for critical states, with different decay exponents depending on the measurement axis. The MINN under $y-$ and $z-$axes measurements decays more slowly than the pre-measurement behavior, but still faster than the slowest decaying correlation prior to measurement, namely the $\left\langle\sigma_{x}^{1}\sigma_{x}^{r}\right\rangle$, which decays with exponent $1/4$ Francesco et al. (2012). For non-critical states, the MINN either becomes constant or decays exponentially. These findings indicate that both NN and MINN can serve as effective probes of the underlying quantum phase transitions in many-body systems, in line with the growing interest in probing quantum phase transitions through quantum information measures.
@@
Figure 3: Results for critical MHC with $L=48$, with a maximum bond dimension $\chi=1000$. (a) Haar-random unitaries are applied in a brick-wall manner, with a measurement rate of $p=0.17$ Zabalo et al. (2020). The measurements within the dotted box are used solely for computing MINN, with all qubits except the first and the $r$-th qubits are projectively measured along the $z$-axis. (b) A comparison between averaged two-point NN $\overline{\mathcal{M}}$ and total nonstabilizerness $\overline{M}$. Each data point is averaged over 3000 trajectories. The inset shows the power-law fit with a constant shift for $\overline{\mathcal{M}}$. (c) The MINN and the average total nonstabilizerness after measurement, with each data point averaged over $4\times 10^{4}$ trajectories. (d) A comparison between the scaling behavior of the MINN and the mutual information $I_{E}$ prior measurement. The decay exponents for $\widetilde{\mathcal{M}}$ and $I_{E}$ are around $0.76$ and $3.31$, respectively.
@@
We then consider the NN and MINN in MHC at a critical measurement rate of $p=0.17$. The circuit structure is shown in Fig. 3(a). We consider the average NN defined as ¯$\overline{\mathcal{M}}=\mathbb{E}_{U}\mathbb{E}_{m}[\mathcal{M}]$, where $\mathbb{E}_{U}$ and $\mathbb{E}_{m}$ represent the averaging over unitary gates and the measurement outcomes, respectively. Since the model is structureless, with only locality as a constraint, $\rho_{1,r}$ does not necessarily have a canonical Pauli basis representation, and the NN must be computed by numerical optimization. As shown in Fig. 3(b), NN is significantly lower than total nonstabilizerness and also exhibits a power-law decay. We then consider the MINN and compare it with the average total nonstabilizerness for post-measurement states, as presented in Fig. 3(c). The measurement direction is insignificant, and we assume it to be along the $z$-axis. It is evident that a greater amount of nonstabilizerness can be removed by local unitary transformations as the distance increases. Furthermore, we observe that the decay exponent of MINN is less than half of that of mutual information prior to measurement, as shown in Fig. 3(d). Since the slowest correlation a state can maintain must decay at an exponent at least half that of mutual information Wolf et al. (2008), this suggests that the measurement induces NN with a longer range than any pre-existing correlations. This phenomenon is akin to entanglement swapping, where measurements induce long-range entanglement between previously uncorrelated qubits. We refer to this effect as “nonstabilizerness swapping”. Moreover, since the entanglement is lower bounded by $\mathcal{M}$ Cao et al. (2024), the presence of nonstabilizerness swapping serves as a sufficient condition for the occurrence of entanglement swapping. It is also worth noting that the occurrence of entanglement swapping serves as a sufficient condition for a state to possess a nontrivial sign structure; that is, the state must exhibit certain negative amplitudes in the measurement basis Lin et al. (2023). Consequently, the phenomenon of nonstabilizerness swapping necessarily implies that the underlying state exhibits a nontrivial sign structure—an alternative diagnostic of a state’s “quantumness” Hastings (2016).
VI Discussions
Our results carry several implications across different fields. First, an exact expression for computing the two-qubit NN is not only valuable for quantifying non-classical resources and classical simulation complexity, but also holds direct relevance to the gravitational back-reaction Cao et al. (2024); Cao (2024). Second, the distinct power-law behavior exhibited by the two-point NN at the critical point provides insight into the scaling of NN in generic conformal field theories. Importantly, NN is invariant under local unitary transformations, indicating that it may be computed within the conformal field theory framework in the continuum limit—analogous to the calculation of entanglement entropy. Our exact results for the two-point case thus serve as a benchmark for such a formalism and we leave it to future work. In contrast, mutual SRE is sensitive to microscopic details, making its continuum description more elusive. Third, nonstabilizerness swapping provides a unified manifestation of the three fundamental quantum resource aspects—entanglement, nonstabilizerness, and sign structure—such that any state exhibiting this phenomenon is necessarily nontrivial from all three perspectives. Identifying further examples of such states remains an intriguing open problem, which we leave for future work.
@@ VI Discussions
It is noteworthy that a related measure, nonstabilizerness entanglement entropy (NsEE), has recently been introduced Qian et al. (2024); Huang et al. (2024). In this formulation, free operations are defined as global Clifford gates instead of local unitary gates, and the computed quantity is the entanglement entropy of the total state rather than its nonstabilizerness. While NsEE and NN aim to characterize the “quantumness” of a system from distinct perspectives, further investigation into the relationship between these two measures would be a valuable direction for future research.
@@ Acknowledgements.
We thank Xiaoyu Dong and Pengfei Zhang for valuable discussions. This work is supported by the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600, No. 24LZ1400100 and No. 2019SHZDZX01, and is sponsored by “Shuguang Program” supported by Shanghai Education Development Foundation and Shanghai Municipal Education Commission.
Appendix A Proof of sub-additivity
In this section, we prove that the NN, defined as $\mathcal{M}_{AB}(\left|\Psi_{AB}\right\rangle)=\min_{U_{A}\otimes U_{B}}M(U_{A}\otimes U_{B}\left|\Psi_{AB}\right\rangle)$, is sub-additive if the nonstabilizerness measure $M$ is additive. By definition, we have:
$\mathcal{M}_{AB}(\left|\Psi_{A_{1}B_{1}}\right\rangle\otimes\left|\Phi_{A_{2}B_{2}}\right\rangle)$ (3)
$=\min_{{U_{A}\otimes U_{B}}}M(\left|\Psi_{A_{1}B_{1}}\right\rangle\otimes\left|\Phi_{A_{2}B_{2}}\right\rangle)$
$=\min_{U_{A}\otimes U_{B}}[M(\left|\Psi_{A_{1}B_{1}}\right\rangle)+M(\left|\Phi_{A_{2}B_{2}}\right\rangle)].$
@@ Appendix A Proof of sub-additivity
$\mathcal{M}_{AB}(\left|\Psi_{A_{1}B_{1}}\right\rangle\otimes\left|\Phi_{A_{2}B_{2}}\right\rangle)$ (4)
$\leq\min_{U_{A_{1}}\otimes U_{B_{1}}}M(\left|\Psi_{A_{1}B_{1}}\right\rangle)+\min_{U_{A_{2}}\otimes U_{B_{2}}}M(\left|\Phi_{A_{2}B_{2}}\right\rangle)$
$=\mathcal{M}_{A_{1}B_{1}}(\left|\Psi_{A_{1}B_{1}}\right\rangle)+\mathcal{M}_{A_{2}B_{2}}(\left|\Phi_{A_{2}B_{2}}\right\rangle).$
@@
Building on this sub-additivity, it is straightforward to establish an upper bound for the NN of a many-body state, provided that it can be brought into product states of two-qubit pure states in canonical form via local unitary transformations. It has been shown that such transformations can be approximately realized using the Petz recovery map for Haar-random states Yoshida and Kitaev (2017). However, the exact determination of such a transformation remains an open problem and is a subject worth further exploration.
Appendix B Proof of theorem
@@
$\rho=\frac{1}{4}\sum_{\mu,\nu}R_{\mu\nu}\sigma_{\mu}\sigma_{\nu},$ (5)
where $\mu,\nu\in\{0,x,y,z\}$, and the coefficients $R$ are determined by $R_{\mu\nu}=\text{Tr}(\rho\sigma_{\mu}\otimes\sigma_{\nu})$ Schlienz and Mahler (1995); Horodecki and Horodecki (1996). The matrix $R$ takes a block structure of the form:
$R=\begin{pmatrix}1&\textbf{a}^{T}\\ \textbf{b}&T\end{pmatrix},$ (6)
where a and b are the Bloch vectors corresponding to the reduced density matrices of qubit $A$ and $B$, respectively, and $T$ represents the correlation matrix. In this representation, the second SRE of the state $\rho$ is concisely expressed as:
$M(\rho)=-\text{ln}\left(\sum_{\mu\nu}{|R_{\mu\nu}|^{4}}\right)-S_{2}(\rho),$ (7)
@@
$\sum_{i}\left|a_{i}\right|^{4}\leq\sum_{i}\left|a_{i}\right|^{2}=\text{const.},$. , (8)
@@
$\sum_{i=1}^{3}\left|T^{\prime}_{i,j}\right|^{4}$ $=\sum_{i=1}^{3}\left|d_{j}O_{A,ij}\right|^{4}=\left|d_{j}\right|^{4}\sum_{i=1}^{3}\left|O_{A,ij}\right|^{4}$ (9)
$\leq\left|d_{j}\right|^{4}\sum_{i=1}^{3}\left|O_{A,ij}\right|^{2}=\left|d_{j}\right|^{4},$
@@
$R=\begin{pmatrix}1&0&0&\text{cos}(2\theta)\\ 0&\text{sin}(2\theta)&0&0\\ 0&0&-\text{sin}(2\theta)&0\\ \text{cos}(2\theta)&0&0&1\end{pmatrix},$ (10)
@@
It is worth noting that while a local unitary transformation always exists that brings a two-qubit pure state into canonical form, this is not generally true for a mixed state. However, a unique normal form exists for any two-qubit mixed state, where the correlation matrix $T$ is diagonal, although the Bloch vectors of the qubits need not be fully polarized F. Verstraete (2002).
Appendix C Probability distribution
In this section, we derive the analytic expression for the probability distribution of NN for the two-qubit Haar random state. It is well known that the Haar random state has an entanglement spectrum $\{\text{cos}(\theta),\text{sin}(\theta)\}$, with the associated probability distribution given by $P(x)=3(1-2x)^{2}$, where we define $x\equiv\text{cos}^{2}(\theta)$ Page (1993); Sen (1996). From the corollary stated in the main text, the NN for a state with such an entanglement spectrum is expressed as:
$\mathcal{M}(x)$ $=\text{ln}\left(\frac{8}{7+\text{cos}(8\theta)}\right)$ (11)
$=\text{ln}\left(\frac{1}{1+4x(x-1)(1-2x)^{2}}\right).$